Operations Report – December 21st

SOL: 11
Name of person filing report: Jesus Meza-Galvan
Non-nominal systems: Crew car.
Notes on non-nominal systems: Crew car sprung a coolant leak that needs to be repaired by mechanic. Mission support is aware and coordinating repairs.
Spirit rover used: No
Hours: (before EVA)
Beginning charge: (Before EVA)
Ending charge: (On return from EVA, before recharging)
Currently charging: Yes
Opportunity rover used: Yes
Hours: 151.0
Beginning charge: 100%
Ending charge: 44 %
Currently charging: No
Curiosity rover used: No
Beginning charge:
Ending charge:
Currently charging: Yes
Perseverance rover used: Yes
Hours: 280.6
Beginning charge: 100%
Ending charge: 31%
Currently charging: No
General notes on rovers: Rovers are functioning properly.
Summary of Hab operations:
WATER USE: 23 gallons
Water (static tank): 266 gallons
Static tank pipe heater (on or off): On
Static tank heater (on or off): On
Toilet tank emptied: Yes.
Summary of internet: Internet is nominal.
Summary of suits and radios: Suits and radios are nominal.
Summary of GreenHab operations:
WATER USE: 2.0 gallons
Heater: On
Supplemental light: On
Parsley – 14 g

Summary of ScienceDome operations: Science dome was used today for geology sample packing. The photolithography set-up in the laminar hood was taken down. The laminar hood is now back to its normal configuration. Photolithography samples and waste material were packed and removed from ScienceDome. Waste will be taken away by the crew when leaving MDRS.
Dual split: Heat or AC: Off
Summary of RAM operations: (Tools used, work done): No work was conducted in the RAM today. Dremel stand used for lithography was returned to the RAM.
Summary of any observatory issues: All systems are nominal.
Summary of health and safety issues: No health or safety issues.
Questions, concerns and requests to Mission Support: None.

Copyright © The Mars Society. All rights reserved. | Main Site